# Exploring Facial Expression Recognition through Semi-Supervised Pretraining and Temporal Modeling

### 18 Mar 2024 v1

@@ Abstract
Facial Expression Recognition (FER) plays a crucial role in computer vision and finds extensive applications across various fields. This paper aims to present our approach for the upcoming 6th Affective Behavior Analysis in-the-Wild (ABAW) competition, scheduled to be held at CVPR2024.. In the facial expression recognition task, The limited size of the FER dataset poses a challenge to the expression recognition model’s generalization ability, resulting in subpar recognition performance. To address this problem, we employ a semi-supervised learning technique to generate expression category pseudo-labels for unlabeled face data. At the same time, we uniformly sampled the labeled facial expression samples and implemented a debiased feedback learning strategy to address the problem of category imbalance in the dataset and the possible data bias in semi-supervised learning. Moreover, , to further compensate for the limitation and bias of features obtained only from static images, we introduced a Temporal Encoder to learn and capture temporal relationships between neighboring expression image features. In the 6th ABAW competition, our method achieved outstanding results on the official validation set, a result that fully confirms the effectiveness and competitiveness of our proposed method.
@@ 1 Introduction
To stimulate interdisciplinary collaboration and address pivotal research inquiries spanning affective computing, machine learning, and multi-modal signal processing, Kollias et al. have spearheaded the Affective Behavior Analysis in-the-wild (ABAW) initiative. The 6th iteration of this workshop and competition is slated to align with the IEEE CVPR conference in 2024[35, 16, 12, 17, 11, 18, 19, 13, 9, 10, 14, 15].
@@
In this study, we propose a two-phase methodology to enhance the recognition and analysis of facial expressions. The first phase, known as the spatial pre-training phase, plays a crucial role in preparing the model for subsequent tasks. During this phase, we leverage the power of semi-supervised learning techniques to generate pseudo-labels for expression categories using unlabeled face data. This approach ensures a sufficiently large training corpus, allowing the model to effectively extract robust facial expression features. To tackle the challenges of category imbalance in the dataset and potential data bias during semi-supervised learning, we adopt two strategies. First, we uniformly sample labeled facial expression instances to address the category imbalance. Additionally, we employ a debiased feedback learning strategy to mitigate the impact of potential data bias. These strategies collectively contribute to training a more robust facial expression recognizer. Moving to the second phase, the temporal learning phase, we aim to further improve the recognition and analysis of facial expressions by capturing the temporal dynamics. In this phase, we freeze the facial expression recognizer trained in the first phase, which has already acquired strong spatial representation capabilities. To incorporate temporal information, we introduce a temporal encoder that learns the temporal relationships between neighboring expression image features. By considering the temporal aspect, we compensate for the inherent feature bias obtained solely from static images. This integration of temporal dynamics enables more accurate and comprehensive dynamic recognition and analysis of facial expressions.
@@
  • To address the problem of scarcity of facial expression data, we applied a semi-supervised learning technique to generate expression category pseudo-labels for unlabeled face data. At the same time, we uniformly sampled the labeled facial expression samples and implemented a debiased feedback learning strategy to solve the problem of category imbalance in the dataset and the possible data bias in the semi-supervised learning, which improves the robustness and generalization ability of the model.
  • In order to compensate for the limitations and biases of features acquired only from static images, we introduce a temporal encoder to learn and capture temporal relationships between neighboring expression image features. This strategy aims to enhance the model’s ability to recognize and analyze the dynamic changes in facial expressions, and achieve more accurate dynamic facial expression recognition.
@@
Figure 1: Framework Description. Our approach is mainly divided into a spatial phase and a temporal phase. (1). The goal of the spatial phase is to expand face expression data by mining large-scale unlabeled faces through a semi-supervised algorithm. (2). The goal of the temporal phase is to do temporal feature enhancement of the image features extracted by the student network in the first phase by means of a temporal encoder, so as to improve the accuracy of recognizing the dynamic facial expressions in the video.
@@ 2.1 Facial Expression Recognition
The facial expression recognition task is a classic task in the field of pattern recognition, and the approach[29, 26, 39, 3, 2, 33] of making full use of fully supervised data once gained very great progress in the field of FER. In recent years, attention has been focused on extending the dataset to obtain larger scale datasets. To solve the label confusion problem between different expression recognition datasets. IPA2LT [37] is the first work to address the annotation inconsistency in different facial expression datasets. They proposed LTNet embedded with a scheme of discovering the latent truth from multiple inconsistent labels and the input images. Ada-CM [22] is the first solution to explore the dynamic confidence margin in Semi-Supervised Deep Facial Expression Recognition. They designed an adaptive confidence margin to dynamically learn on all unlabeled data for the model’s training and conducted a feature-level contrastive objective to learn effective features by applying the InfoNCE[30] loss. Face2Exp [36] proposed the Meta-Face2Exp framework to extract de-biased knowledge from auxiliary FR data through the meta optimization framework. utilized unlabeled face data to enhance expression recognition through the meta optimization framework.
2.2 Learning with Unlabeled Data
An important direction for learning methods using unlabeled data is semi-supervised learning. A popular class of Semi-Supervised Learning methods is to generate an artificial label for an unlabeled image and train the model to predict that artificial label on the input of an unlabeled image [21, 31]. Similarly, consistent regularization [1, 20, 25] uses the predictive distribution of the model to obtain an artificial label after randomly modifying the input or model function. An artificial label is obtained using the predictive distribution of the model after randomly modifying the input or the model function. Fixmatch[27] combines the advantages of these two methods, which use weak data augmentation and strong data augmentation of samples to achieve consistent regularization and obtain pseudo-labeled data by samples with confidence levels greater than a threshold. But the problem of this approach is that his modeling ability is limited in the early stage of model training because the threshold value is fixed. To solve this problem, Flexmatch[38] proposed Curriculum Pseudo Labeling (CPL), a curriculum learning approach to leverage unlabeled data according to the model’s learning status. It can flexibly adjust thresholds for different classes at each time step to let pass informative unlabeled data and their pseudo labels. In addition, Dash[32] performs selection by retaining only samples with losses less than a given threshold in each update iteration, which is dynamically adjusted through iterations.
@@ 3.1 Overview
As shown in Figure.1, our approach uses multiple techniques to improve facial expression recognition (FER) in two phases. The first phase, the spatial phase, uses semi-supervised learning to increase the FER dataset by using facial recognition (FR) data. This helps in training a better image feature extractor. We use two neural networks, a teacher and a student, with the same structure but different weights.In this phase, we improve the networks by using a debiasing method. This involves comparing biased FR data (unlabeled) with debiased FER data (labeled) to create pseudo-labels. The teacher network learns from balanced FER data and makes these pseudo-labels, which the student network then uses to learn from the FR data, adjusting based on feedback to improve accuracy.In the training, the teacher and student networks update alternately with the debiasing method to better process FR data, improving accuracy and debiasing. The second phase, the temporal phase, uses the student network (fixed after the first phase) to extract image features. To address biases from static images, we use a temporal encoder to understand time-related feature relationships, improving system accuracy. This enables dynamic recognition and analysis of facial expressions in videos, with a classifier providing frame-by-frame predictions.
@@ 3.3.2 Teacher Network
The $\mathrm{Aug(\cdot)}$ denotes data augmentation, this approach utilizes intensive data augmentation techniques, including rotation, removal, and pixel-level image manipulation.
@@ 3.3.3 Debiasing Mechanism
The behavior of new adaptation network on balanced FER data is used as the measurement. Specifically, when adaptation network on FR data and new adaptation network on FER data have the same cognitive (i.e., sign of their gradients), we obtain positive feedback coeffi- cients (i.e., positive value for Lf ), which encourages the up- date of base network by using the current direction of gradients. When adaptation network on FR data and new adaptation network on FER data have different cognitive (i.e., sign of their gradients), feedback coefficients will give a negative sign, which punishes the update of base network by using the adverse direction of current gradients. In this way, feedback is used as the reward signal that goes back through the base network, determining how parameters of the base network impact the gradient of the adaptation network for debiased feature extraction. As a result, the learning of adaptation network on biased FR data will perform de-biased behavior that is consistent with the evaluation procedure on balanced FER data.
@@ 3.5 Post-Process
Table 1: Ablation study results on the official validation set.
@@
Baseline $\checkmark$ 27.65
SSL $\checkmark$ $\checkmark$ 40.96
SSL $\checkmark$ $\checkmark$ 41.57
SSL + Temporal $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 43.77
SSL + Temporal+ Post-process $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 45.43
@@ 4.1 Datasets
FER Datasets. The 6th Workshop and Competition on Affective Behavior Analysis in-the-wild introduces the Aff-wild2 database as its cornerstone dataset, a compilation that encompasses substantial contributions from a series of studies [35, 16, 12, 17, 11, 18, 19, 13, 9, 10, 14, 15]. In the context of the EXPR Classification Challenge, the Aff-wild2 database, which is audiovisual in nature, comprises approximately 548 videos, culminating in around 2.7 million frames. These frames are meticulously annotated to reflect the six fundamental expressions—anger, disgust, fear, happiness, sadness, surprise—alongside a neutral state and an additional ’other’ category to denote expressions or affective states beyond the basic ones.
Furthermore, our research incorporates data from external facial expression databases, notably AffectNet[24] and ExpW[40], to enrich our analysis. AffectNet, which aggregates around 1 million facial images sourced from the internet, assigns eleven distinct emotion and non-emotion categorical labels. For our purposes, we utilized images corresponding to the first seven categories. Meanwhile, the Expression in-the-Wild Database (ExpW) amasses 91,793 facial images procured via Google image search, with each image manually classified into one of the seven foundational expression categories.
In this study, we achieved a balanced dataset by uniformly extracting 8,000 labeled images for each category from the Aff-Wild2 collection. To further enhance the diversity of our dataset and bolster the generalizability of our model, we also randomly selected 8,000 images for each category from a merged dataset comprising AffecNet and ExpW. For images falling into the ’other’ category, we relied exclusively on the Aff-Wild2 dataset, given the absence of such images in the alternate databases. The compilation is rounded off with the inclusion of the remaining images from these datasets as unlabeled samples, serving to augment our dataset’s comprehensiveness.
@@ 4.4 Results
4.5 Ablation Study
To validate the efficacy of our approach, we executed ablation studies on each component and strategy within our method, with the outcomes presented in Table.LABEL:table1 It is evident that the application of the SSL (Semi-Supervised Learning) technique significantly enhances the recognition performance by 13.31%. Further augmentation of the facial expression dataset can lead to an increase in accuracy. Moreover, incorporating the temporal encoder results in an additional 2.2% improvement in accuracy, underscoring the importance of temporal learning. Ultimately, through post-processing, the model achieves an accuracy rate of 45.43%.

### 19 Mar 2024 v2

@@ Abstract
Facial Expression Recognition (FER) plays a crucial role in computer vision and finds extensive applications across various fields. This paper aims to present our approach for the upcoming 6th Affective Behavior Analysis in-the-Wild (ABAW) competition, scheduled to be held at CVPR2024. In the facial expression recognition task, The limited size of the FER dataset poses a challenge to the expression recognition model’s generalization ability, resulting in subpar recognition performance. To address this problem, we employ a semi-supervised learning technique to generate expression category pseudo-labels for unlabeled face data. At the same time, we uniformly sampled the labeled facial expression samples and implemented a debiased feedback learning strategy to address the problem of category imbalance in the dataset and the possible data bias in semi-supervised learning. Moreover, to further compensate for the limitation and bias of features obtained only from static images, we introduced a Temporal Encoder to learn and capture temporal relationships between neighbouring expression image features. In the 6th ABAW competition, our method achieved outstanding results on the official validation set, a result that fully confirms the effectiveness and competitiveness of our proposed method.
@@ 1 Introduction
To stimulate interdisciplinary collaboration and address pivotal research inquiries spanning affective computing, machine learning, and multi-modal signal processing, Kollias et al. have spearheaded the Affective Behavior Analysis in-the-wild (ABAW) initiative. The 6th ABAW workshop and competition is slated to align with the IEEE CVPR conference in 2024[35, 16, 12, 17, 11, 18, 19, 13, 9, 10, 14, 15].
@@
In this study, we propose a two-phase methodology to enhance the recognition and analysis of facial expressions. The first phase, known as the spatial pre-training phase, plays a crucial role in preparing the model for subsequent tasks. This phase is an improvement on the method[36]. During this phase, we leverage the power of semi-supervised learning techniques to generate pseudo-labels for expression categories using unlabeled face data. This approach ensures a sufficiently large training corpus, allowing the model to effectively extract robust facial expression features. To tackle the challenges of category imbalance in the dataset and potential data bias during semi-supervised learning, we adopt two strategies. First, we uniformly sample labeled facial expression instances to address the category imbalance. Additionally, we employ a debiased feedback learning strategy to mitigate the impact of potential data bias. These strategies collectively contribute to training a more robust facial expression recognizer. Moving to the second phase, the Temporal Refine phase, we aim to further improve the recognition and analysis of facial expressions by capturing the temporal dynamics. In this phase, we freeze the facial expression recognizer trained in the first phase, which has already acquired strong spatial representation capabilities. To incorporate temporal information, we introduce a temporal encoder that learns the temporal relationships between neighbouring expression image features. By considering the temporal aspect, we compensate for the inherent feature bias obtained solely from static images. This integration of temporal dynamics enables more accurate and comprehensive dynamic recognition and analysis of facial expressions.
@@
  • To address the problem of scarcity of facial expression data, we applied a semi-supervised learning technique to generate expression category pseudo-labels for unlabeled face data. At the same time, we uniformly sampled the labeled facial expression samples and implemented a debiased feedback learning strategy to solve the problem of category imbalance in the dataset and the possible data bias in semi-supervised learning.
  • In order to compensate for the limitations and biases of features acquired only from static images, we introduce a temporal encoder to learn and capture temporal relationships between neighbouring expression image features. This strategy aims to enhance the model’s ability to recognize and analyze the dynamic changes in facial expressions and achieve more accurate dynamic facial expression recognition.
@@
Figure 1: Framework Description. Our approach is mainly divided into a Spatial Pretraining phase and a Temporal Refine phase. (1). The goal of the Spatial Pretraining phase is to expand face expression data by mining large-scale unlabeled faces through a semi-supervised algorithm. (2). The goal of the Temporal Refine phase is to do temporal feature enhancement of the image features extracted by the student network in the first phase by means of a temporal encoder, so as to improve the accuracy of recognizing the dynamic facial expressions in the video.
@@ 2.1 Facial Expression Recognition
The task of recognizing facial expressions is a foundational challenge in pattern recognition. Techniques that leverage fully supervised data have significantly advanced the field of Facial Expression Recognition (FER), as evidenced by groundbreaking research[29, 26, 40, 3, 33]. However, datasets for facial expressions face notable limitations, including a lack of diversity and significant category imbalances. In response, recent efforts have shifted towards expanding these datasets to enrich the variety of facial expressions available for analysis. One pioneering approach to address the challenge of inconsistent labeling across different facial expression datasets is the IPA2LT framework [37]. This methodology introduced the LTNet scheme, an innovative strategy for uncovering the underlying truths among diverse, inconsistent labels through the use of embedded analysis.In the realm of semi-supervised learning for FER, Ada-CM[22] emerged as the first to investigate the concept of dynamic confidence. By designing an adaptive confidence margin, this approach innovatively adapts during training to maximize learning from unlabeled data through feature-level comparisons, utilizing the InfoNCE loss[30] to capture valuable features effectively.Further advancing the field, Face2Exp[36] introduced the Meta-Face2Exp framework. This methodology utilizes a meta-optimization framework to derive unbiased knowledge from auxiliary Facial Recognition (FR) data, showcasing a novel approach to knowledge extraction. Lastly, an innovative technique proposed by Zhang et al.[39] employs rebalancing attention mapping to regularize models. This allows for the extraction of transformation-invariant information from secondary categories across all training samples, addressing the critical issue of data imbalance by focusing on underrepresented categories.
2.2 Semi-Supervised Learning
Semi-supervised learning, focusing on unlabeled data, is a vital area within learning algorithms. A prominent approach in Semi-Supervised Learning is generating synthetic labels for unlabeled images and guiding the algorithm to identify these labels for such image inputs, as discussed in the works [21, 31]. Similarly, consistent regularization, according to [1, 20, 25] utilizes the model’s predictive capacity to form a synthetic label by altering the input or model’s operations randomly. Fixmatch[27] integrates the benefits of generating synthetic labels and applying both light and intensive data augmentations to ensure consistent regularization and generate pseudo-labeled data from samples exceeding a confidence threshold. However, its limitation arises early in model training due to the static threshold.To mitigate this, Flexmatch[38] proposes Curriculum Pseudo Labeling (CPL), leveraging unlabeled data based on the model’s learning stage and dynamically adjusting thresholds for various classes, thus allowing the inclusion of informative unlabeled data and their pseudo labels. Additionally, Dash [32] employs a dynamic threshold to keep only the samples with losses under a specific threshold during updates.
@@ 3.1 Overview
As shown in Figure.1, our approach uses multiple techniques to improve facial expression recognition (FER) in two phases. The first phase, the Spatial Pretraining phase, uses semi-supervised learning to increase the FER dataset by using facial recognition (FR) data. This helps in training a better image feature extractor. We use two neural networks, a teacher and a student, with the same structure but different weights. In this phase, we improve the networks by using a debiasing method. This involves comparing biased FR data (unlabeled) with debiased FER data (labeled) to create pseudo-labels. The teacher network learns from balanced FER data and makes these pseudo-labels, which the student network then uses to learn from the FR data, adjusting based on feedback to improve accuracy. In the training, the teacher and student networks update alternately with the debiasing method to better process FR data, improving accuracy and debiasing. The second phase, the Temporal Refine phase, uses the student network (fixed after the first phase) to extract image features. To address biases from static images, we use a temporal encoder to understand time-related feature relationships, improving system accuracy. This enables dynamic recognition and analysis of facial expressions in videos, with a classifier providing frame-by-frame predictions.
@@ 3.3.2 Teacher Network
The $\mathrm{Aug(\cdot)}$ denotes strong data augmentation, this approach utilizes intensive data augmentation techniques, including rotation, removal, and pixel-level image manipulation.
@@ 3.3.3 Debiasing Mechanism
The performance of a novel student network on balanced Facial Expression Recognition (FER) data serves as the criterion for evaluation. More precisely, positive feedback coefficients are achieved (indicating a favourable $\mathcal{L}_{f}$ value) when both the student network operating on Face Recognition (FR) data and the novel student network focused on FER data share identical gradient orientations. This scenario promotes the modification of the teacher network through the employment of the gradient’s current trajectory. Conversely, should the student network concerning FR data and the novel student network for FER data exhibit contrasting gradient orientations, the feedback coefficients assume a negative value. This acts as a deterrent to the teacher network’s adjustment, advocating for the use of an opposing gradient direction. Thus, feedback operates as a crucial reward signal, retroactively influencing the teacher network by dictating the influence of its parameters on the student network’s gradient for the extraction of unbiased features.
@@ 3.5 Post-Process
Table 1: Ablation study results on the validation set.
@@
Baseline $\checkmark$ 23.00
SSL $\checkmark$ $\checkmark$ 39.96
SSL $\checkmark$ $\checkmark$ 40.57
SSL + Temporal $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 42.77
SSL + Temporal+ Post-process $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 44.43
@@ 4.1 Datasets
FER Datasets.
The 6th Workshop and Competition on Affective Behavior Analysis in-the-wild unveiled the Aff-wild2 database, a pivotal collection drawn from a range of studies[35, 16, 12, 17, 11, 18, 19, 13, 9, 10, 14, 15]. This resource was central to the EXPR Classification Challenge, featuring an audio-visual dataset comprised of 548 videos and approximately 2.7 million frames annotated for six fundamental facial expressions, as well as neutral and ’other’ categories. To enrich the dataset’s depth, we incorporated data from the AffectNet and ExpW databases; AffectNet contributed around 1 million facial images categorized into 11 emotions, while ExpW offered 91,793 images across seven expression categories. This integration aimed to enhance the comprehensiveness and utility of our analysis.
To further enhance the diversity of our dataset and bolster the generalizability of our model, we also randomly selected 8,000 images for each category from a merged dataset comprising AffecNet and ExpW. For images falling into the ’other’ category, we relied exclusively on the Aff-Wild2 dataset, given the absence of such images in the alternate databases. The compilation is rounded off with the inclusion of the remaining images from these datasets as unlabeled samples, serving to augment our dataset’s comprehensiveness.
@@ 4.4 Results
To validate the efficacy of our approach, we executed ablation studies on each component and strategy within our method, with the outcomes presented in Table.1 It is evident that the application of the SSL (Semi-Supervised Learning) technique significantly enhances the recognition performance by 16.96%. Further augmentation of the facial expression dataset can lead to an increase in accuracy. Moreover, incorporating the temporal encoder results in an additional 2.2% improvement in accuracy, underscoring the importance of temporal learning. Ultimately, through post-processing, the model achieves an accuracy rate of 44.43%.
@@ 5 Conclusion
In this paper, we propose a two-phase approach to improve facial expression recognition. The first phase is called the spatial pre-training phase, in which, in order to address the problem of scarcity of facial expression data, we employ a semi-supervised learning technique to generate expression category pseudo-labels for unlabeled facial data. At the same time, we uniformly sampled the labeled facial expression samples and implemented a debiased feedback learning strategy to address the problem of category imbalance in the dataset and possible data bias in semi-supervised learning. The second phase is the temporal refine phase. In this phase, to compensate for the limitation and bias of obtaining features only from static images, we introduced a temporal encoder to learn and capture the temporal relationship between the features of neighbouring expression images to achieve more accurate dynamic facial expression recognition.